Project

General

Profile

Missing Channel Table

Currently (last update 1/11/2011)there are many missing channels at the NDOS detector. All of them were found by running the DCMs in the DSO mode readout (Pedestal data).
The runs were taken with 400V and the RMS of the APD's current for individual pixels was calculated.

The criteria for choosing a good pixel and FEB was:

  • A "good pixel" has an RMS between 4.0 and 100.0 (these two numbers were obtained empirically, by eye). If it is less than 4.0, it is too quiet and doesn't react to the input voltage. If it is more than 100.0, then the pixel is considered too noisy. Typically those quiet pixels had an RMS of 0.0-2.0 and those noisy pixels had an RMS of 200-1400.
  • A "good FEB" has at least 4 good pixels. Some FEBs had 2 or 3 "good pixels", but we drop them out.

Currently all bad pixels or bad FEBs are masked off in the DAQ.

In this table the missing channels are shown together with the initial problem and current hardware status:

DCM FEB Number Initial Problem Current Status
1-1 0 Bad pixels 7
1 Bad FEB
2 Software problem (wouldn't program DCM with this FEB)
3 Bad pixels 14 26
4 Bad FEB
5 Bad FEB
8 Bad pixels 31
9 Bad pixels 1 6 7 9 15 22 30 31
10 Bad pixels 30
11 Bad FEB
12 Bad pixels 8 9 10 12 14 15 17 22 24 26 27 28 29 30 31
13 Bad pixels 6 7 10 11 12 13 14 15 16 23 26 28 29 30 31
14 Bad FEB
15 Bad pixels 0 1 8 9 25 26 27
16 Bad FEB Removed APD (SN H32) from Block 0 Plane 0 Position 1 on Jan 1, 2011. Observed to be generally dirty.
17 Bad pixels 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
18 Bad FEB Removed APD (SN H60) from Block 0 Plane 4 Position 1 on Jan 1, 2011. Observed to be generally dirty.
20 Bad FEB
21 Bad pixels 23
22 Bad pixels 25
23 Bad pixels 0 1 5 8 10 11 12 13 14 15 17 18 19 21 22 23 28 29 30 31
24 Bad FEB
26 Bad FEB
27 Bad FEB
28 Bad FEB
29 Bad FEB
30 Bad FEB APD (SN H051) Removed from detector 12/15/2010 (Some debris on APD surface).
32 Bad FEB
33 Bad FEB
34 Bad pixels 1 6 14 15 16 24 26 27 28 29 30 31
35 Bad FEB
36 Bad pixels 15 16
37 Bad FEB
38 Bad FEB
39 Bad FEB
42 Bad pixels 8 24 25 26 27 28 29 30 31
48 DCM firmware issue
50 Bad pixels 23
51 Bad FEB
52 Bad FEB
56 Bad pixels 8 9 11
57 Bad pixels 6 15
58 DCM firmware issue
59 Bad pixels 26 27 28 29 30
60 Bad pixels 1 4 5 6 7 8 9 10 11 12 13 14 15 16 17 19 20 21 23 27
61 Bad pixels 13
1-2 2 Software problem (wouldn't program DCM with this FEB)
3 Bad FEB
13 Bad FEB
14 Bad pixels 0 1 2 3 5
15 Bad pixels 23 24 25 26 27 28 29 30 31
20 Bad FEB
30 Bad FEB
33 Bad pixels 16 25
37 Bad pixels 8 9 10 12 13 14 15 17 24 25 26 27
38 Bad pixels 14
39 Bad pixels 2 4 7 8 9 10 11 12 13 14 15 16 17 19 22 24 25 26 28
40 Bad pixels 16 17
41 Bad pixels 6 8 10 11 13 15 16 24 25 26 27 28 29 30 31
42 Bad pixels 26 28
43 Bad pixels 24 25
45 Bad pixels 9 10 13 22 24 25 26 27 29 30
48 DCM firmware issue
58 Bad FEB
61 Bad pixels 19
1-3 2 Software problem (wouldn't program DCM with this FEB)
5 Bad pixels 4 15 16 25
32 Bad pixels 1 14 15
33 Bad pixels 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
34 Bad pixels 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 29 30 31
40 Bad pixels 13
42 Bad pixels 9
46 Bad FEB
2-1 2 Software problem (wouldn't program DCM with this FEB)
3 Bad FEB
4 Bad FEB
7 Bad pixels 8 9 22 23 30
8 Bad pixels 8 11 22 23 30 31
10 Bad pixels 6 15 16 24 25 27
11 Bad pixels 22
12 Bad pixels 20 22 24 26 27 29 30 31
13 Bad pixels 22 24 27 28 30
14 Bad pixels 12 13
15 Bad pixels 29
19 Bad pixels 14 15 16 24 25
20 Bad FEB
21 Bad pixels 17 18 19 24 26 27
22 Bad FEB
23 Bad FEB
24 Bad FEB
26 Bad pixels 22 23
28 Bad pixels 24 25
30 Bad FEB
32 Bad pixels 6 7 24 25
33 Bad pixels 6 8 9 10 11 12 13 14 15 16 20 22 24 25 28 30 31
34 Bad FEB
37 Bad pixels 10
40 Bad FEB
41 Bad FEB
42 Bad pixels 10
44 Bad pixels 22
48 DCM firmware issue
49 Bad pixels 14 15 22 23 24 25 26 28
52 Bad FEB
53 Bad FEB
54 Bad pixels 6 14 15
55 Bad FEB
56 Bad pixels 2 3 4 5
57 Bad FEB
58 DCM firmware issue
60 Bad FEB
61 Bad pixels 5
62 Bad FEB
2-2 0 Bad pixels 1 9 30
1 Bad FEB
2 Software problem (wouldn't program DCM with this FEB)
3 Bad pixels 22
4 Bad pixels 14 15 16 17 24 25 26
6 Bad pixels 14 15 16 17 19 24 25
7 Bad pixels 1 22 24 25 26 27 28 29 30 31
8 Bad pixels 12 15 16 24 25 28 30 31
9 Bad FEB
10 Bad FEB APD (SN G165) removed from detector 12/17/2010 (Fiber marks visible on APD)
11 Bad pixels 0 1 2 3 4 5 6 7 8 9 10 11 12 17 18 21 22 23 24 25 26 28 29 30 31 APD (SN G163) removed from detector 12/17/2010 (Discoloration on APD). Shipped to UMn 12/22/2010
12 Bad FEB
13 Bad pixels 14
15 Bad pixels 24 25 27
16 Bad pixels 16
17 Bad pixels 18
18 Bad pixels 23 24 25 26 27 28 29 30 31
19 Bad pixels 16 17 19
20 Bad FEB
21 Bad FEB
22 Bad pixels 10 22 23
23 Bad pixels 23
25 Bad pixels 20 21 25
26 Bad pixels 10 13 21 22 24 25 26 27 28 29 30 31
27 Bad pixels 24 25
30 Bad FEB
32 Bad FEB
33 Bad pixels 8
34 Bad pixels 22 26 30 31
36 Bad FEB Removed APD (SN G031) from Block 3 Plane 10 Position 2 on Jan 1, 2011. Oily.
37 Bad FEB Removed APD (SN G032) from Block 3 Plane 12 Position 2 on Jan 1, 2011. Oily.
38 Bad FEB Removed APD (SN G037) from Block 3 Plane 14 Position 2 on Jan 1, 2011. Oily.
39 Bad FEB Removed APD (SN G038) from Block 3 Plane 16 Position 2 on Jan 1, 2011. Oily.
40 Bad FEB
41 Bad pixels 16
42 Bad FEB
43 Bad FEB
44 Bad FEB
45 Bad FEB
46 Bad FEB
48 DCM firmware issue Removed APD (SN G166) from Block 3 Plane 2 Position 1 on Jan 1, 2011. Observed to be generally dirty.
49 Bad FEB Removed APD (SN G219) from Block 3 Plane 4 Position 1 on Jan 1, 2011. Observed fiber contact marks.
50 Bad FEB Removed APD (SN G211) from Block 3 Plane 6 Position 1 on Jan 1, 2011. Observed to be generally dirty.
51 Bad FEB Removed APD (SN G180) from Block 3 Plane 8 Position 1 on Jan 1, 2011. Oily.
52 Bad pixels 12
53 Bad FEB
54 Bad FEB
56 Bad FEB Removed APD (SN G179) from detector 12/17/2010 (Oily). Shipped to UMn on 12/22/2010. A new APD (SN H19) was installed after cleaning the optical connector 01/05/2011. Removed APD (SN H19) from Block 3 Plane 18 Position 1 on Jan 1, 2011. Oily and Fiber marks.
57 Bad FEB Removed APD (SN G222) from detector 12/17/2010 (Oily). Shipped to UMn on 12/22/2010. A new APD (SN H18) was installed after cleaning the optical connector 01/05/2011.
58 DCM firmware issue
59 Bad FEB APD (SN H059) was cleaned of oil and reinstalled 01/05/2011
60 Bad FEB
61 Bad FEB
62 Bad pixels 0 1 2 3 4 5 6 8 9 10 11 12 13 15 16 17 18 19 20 22 23 26 27 28 29 31
2-3 2 Software problem (wouldn't program DCM with this FEB)
4 Bad FEB
11 Bad pixels 12
12 Bad pixels 0 2 3 8 9 15 25 26 27 28 29 30 31
13 Bad pixels 1 13 24 27
32 Bad pixels 1 3 5 7 9 11 13 15 16 17 18 19 20 21 22 23 24 25 26 27 28 31
33 Bad pixels 16 24 25 27
35 Bad pixels 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31
36 Bad pixels 0 1 2 3 4 5 8 9 10 11 13 14
37 Bad pixels 1
38 Bad FEB
39 Bad FEB APD (SN H028) removed from detector 12/17/2010 (Scratches)
40 Bad pixels 0 1 2 3 4 5 6 8 14 15
41 Bad FEB
43 Bad FEB